OpenStudy (anonymous):

Use the elimination method to solve the system of equations. Choose the correct ordered pair. -x + 5y = -4 4x + 3y = 16 A. (19, 3) B. (14, 2) C. (4, 0) D. (9, 1)

OpenStudy (shinalcantara):

By elimination method: -x + 5y = -4 -----equation 1 4x + 3y = 16 -----equation 2 multiply equation 1 with 4 4(-x + 5y = -4) -4x + 20y = -16 ----equation 3 add equations 2 and 3 4x + 3y = 16 -4x + 20y = -16 0 + 23y = 0 y = 0 substitute the value of y to equation 1 -x + 5y = -4 -x + 5(0) = -4 -x = -4 x = 4

OpenStudy (anonymous):

Thanks ^^

OpenStudy (shinalcantara):

yw :)